# QCDSF Collaboration

### v1

@@ III Feynman-Hellmann method
Table 1: Lattice parameters of the gauge ensembles used in this work Bietenholz et al. (2011), including $m_{p}$ Batelaan (2023), $Z_{V}$ Bickerton (2020), $g_{S}$ Smail et al. (2023) and $\langle x\rangle$ Dragos et al. (2016).
$\begin{array}[]{c c c c c c c c c c c c c}N_{f}&N_{L}^{3}\times N_{T}&a\,[\text{fm}]&\beta&\kappa_{\ell},\kappa_{s}&am_{p}&am_{0}&m_{\pi}\,[\text{MeV}]&Z_{V}&g^{u}_{S}&g^{d}_{S}&\langle x\rangle_{u}&\langle x\rangle_{d}\\ \hline\cr\hline\cr 2+1&32^{3}\times 64&0.074&5.50&0.120900&0.4646(43)&0.00674(7)&470&0.85&4.26(10)&2.77(7)&-&-\\ 2+1&48^{3}\times 96&0.068&5.65&0.122005&0.3860(44)&0.00545(15)&410&0.86&4.83(23)&3.15(14)&0.403(1)&0.173(5)\\ 2+1&48^{3}\times 96&0.058&5.80&0.122810&0.3552(27)&0.00574(5)&430&0.88&4.47(28)&2.90(17)&-&-\\ \end{array}$
We perform our calculation of the proton subtraction function on three gauge ensembles generated by the QCDSF collaboration Bietenholz et al. (2011). All ensembles have $2+1$ quark flavours at the SU(3) flavour symmetric point, and have an unphysical pion mass—approximately three times the physical pion mass. In expectation of the presence of large discretisation artefacts, we utilise three ensembles with relatively fine values of the lattice spacing: $a=0.074,0.068,0.058\;\text{fm}$. See Table 1 for more details.
For each gauge ensemble, we determine the subtraction function for a set of $Q^{2}$ values given in Table 2. The overall range of kinematics is $1\lesssim Q^{2}\lesssim 16\;\text{GeV}^{2}$, which pushes below the threshold of the perturbative region, and up to the region where we expect the only significant contributions to be from the OPE, Eq. (7). To determine the subtraction function we extract the energy shifts for each $\mathbf{q}$ following the methodology given in Section III for two values of the external coupling $\lambda\in[0.0125,0.0250]$ as has been done in the same way in previous Feynman-Hellmann calculations of the Compton amplitude Can et al. (2020); Batelaan et al. (2023). A single plateau fit is performed to each ratio given in Eq. (14) with the plateau regions chosen following a covariance-matrix based $\chi^{2}$ analysis. This allows us to perform simple quadratic fits in $\lambda$ to determine the subtraction functions following Eqs. (11) and (15).
Table 2: Kinematics for the all the results in this work.
@@ V.1 Evaluation of the discretisation artefact
In the bottom panel of Fig. 6, we plot the matrix element, $Z_{V}^{2}\langle p|\bar{\psi}\psi|p\rangle$, against the lattice spacing $a$. We see that there is a slight linear dependence of the matrix element on the lattice spacing $a$. Thus we anticipate a slightly larger correction to the lattice results at a larger lattice spacing, as is suggested by Fig. 3. Moreover, the ‘jitters’ we see in the Wilson coefficient largely explain the small fluctuations with $Q^{2}$ in our lattice subtraction function.
@@ V.2 Improved subtraction function
Finally, we determine the improved subtraction function by removing the $\mathcal{O}(a)$ discretisation artefacts calculated in Section IV using Eq. (18). Combining the direct lattice subtraction function, Fig. 3, with the leading discretisation artefacts, Fig. 6, we present our final results for the up ($uu$) and down ($dd$) contributions in Fig. 7, and for the proton in Fig. 8. Included in both plots (dotted curve) is the continuum OPE calculated from Eq. (7).
@@ VI Summary and Conclusions
By applying the Feynman-Hellmann method to the forward Compton amplitude, we have presented a determination of the subtraction function from first-principles. Furthermore, by performing a LOPE of the forward Compton amplitude, we successfully quantified and removed leading order Wilson discretisation artefacts, thereby placing the raw lattice results in agreement with predictions. In doing so, we demonstrated that the subtraction function indeed follows an asymptotic $Q^{-2}$ behaviour both at large $Q^{2}$ as well as intermediate momenta. Furthermore, the result indicates a smooth transition between the low and high $Q^{2}$ regimes. The results presented here established a pathway for improved lattice calculations of the subtraction function to deliver improved estimates of the proton-neutron mass difference and the proton charge radius.

### v2

@@ III Feynman-Hellmann method
Table 1: Lattice parameters and other relevant quantities of the gauge ensembles used in this work Bietenholz et al. (2011), including $m_{p}$ Batelaan (2023), $Z_{V}$ Bickerton (2020), $g_{S}$ Smail et al. (2023) and $\langle x\rangle$ Dragos et al. (2016).
$\begin{array}[]{c c c c c c c c c c c c c}N_{L}^{3}\times N_{T}&a\,[\text{fm}]&\beta&\kappa_{\ell},\kappa_{s}&am_{p}&am_{0}&m_{\pi}\,[\text{MeV}]&m_{\pi}L&Z_{V}&g^{u}_{S}&g^{d}_{S}&\langle x\rangle_{u}&\langle x\rangle_{d}\\ \hline\cr\hline\cr 32^{3}\times 64&0.074&5.50&0.120900&0.4646(43)&0.00674(7)&470&5.6&0.85&4.26(10)&2.77(7)&-&-\\ 48^{3}\times 96&0.068&5.65&0.122005&0.3860(44)&0.00545(15)&410&6.8&0.86&4.83(23)&3.15(14)&0.403(1)&0.173(5)\\ 48^{3}\times 96&0.058&5.80&0.122810&0.3552(27)&0.00574(5)&430&6.1&0.88&4.47(28)&2.90(17)&-&-\\ \end{array}$
We perform our calculation of the proton subtraction function on three gauge ensembles generated by the QCDSF collaboration Bietenholz et al. (2011) using a non-perturbative, $\mathcal{O}(a)$-improved clover fermion action with unimproved current operators Cundy et al. (2009). All ensembles have $2+1$ quark flavours at the SU(3) flavour symmetric point, and have an unphysical pion mass—approximately three times the physical pion mass. In expectation of the presence of large discretisation artefacts, we utilise three ensembles with relatively fine values of the lattice spacing: $a=0.074,0.068,0.058\;\text{fm}$. See Table 1 for more details.
For each gauge ensemble, we determine the subtraction function for a set of $Q^{2}$ values given in Table 2, which are restricted by our Feynman-Hellmann procedure to $q_{4}=0$. For ease in isolating the subtraction function, we have also introduced the kinematic constraint $q_{3}=0$. The overall range of kinematics is $1\lesssim Q^{2}\lesssim 16\;\text{GeV}^{2}$, which pushes below the threshold of the perturbative region, and up to the region where we expect the only significant contributions to be from the OPE, Eq. (7). To determine the subtraction function we extract the energy shifts for each 3-momentum $\mathbf{q}$ of the external current, following the methodology given in Section III for two values of the external coupling $\lambda\in[0.0125,0.0250]$ as has been done in the same way in previous Feynman-Hellmann calculations of the Compton amplitude Can et al. (2020); Batelaan et al. (2023). A single plateau fit is performed to each ratio given in Eq. (14) with the plateau regions chosen following a covariance-matrix based $\chi^{2}$ analysis. This allows us to perform simple quadratic fits in $\lambda$ to determine the subtraction functions following Eqs. (11) and (15).
Table 2: Momentum transfers of the external current.
@@ V.1 Evaluation of the discretisation artefact
In the bottom panel of Fig. 6, we plot the matrix element, $Z_{V}^{2}\langle p|\bar{\psi}\psi|p\rangle$, against the lattice spacing $a$. We see that there is a linear dependence of the matrix element on the lattice spacing $a$. Thus we anticipate a slightly larger correction to the lattice results at a larger lattice spacing, as is suggested by Fig. 3. Moreover, the ‘jitters’ we see in the Wilson coefficient largely explain the small fluctuations with $Q^{2}$ in our lattice subtraction function.
@@ V.2 Improved subtraction function
Finally, we determine the improved subtraction function by removing the $\mathcal{O}(a)$ discretisation artefacts calculated in Section IV using Eq. (18). Combining the direct lattice subtraction function, Fig. 3, with the leading discretisation artefacts, Fig. 6, we present our final results for the up ($uu$) and down ($dd$) contributions in Fig. 7, and for the proton in Fig. 8. Included in both plots (dotted curve) is the continuum OPE calculated from Eq. (7) using $g_{S}$ and $\langle x\rangle$ from Table 1.The results presented in Fig. 8 are listed in Table 3 of Appendix B.
@@
As a consequence of the different subtraction point used in Ref. Fu et al. (2025), we do not provide a direct comparison with the two results here, but we intend to pursue such a comparison in future work.
@@ VI Summary and Conclusions
By applying the Feynman-Hellmann method to the forward Compton amplitude, we have presented a determination of the subtraction function from first-principles. Furthermore, by performing a LOPE of the forward Compton amplitude, we successfully quantified and removed leading order Wilson discretisation artefacts, thereby placing the raw lattice results in agreement with predictions. In doing so, we demonstrated that the subtraction function indeed follows an asymptotic $Q^{-2}$ behaviour both at large $Q^{2}$ as well as intermediate momenta. Furthermore, the result indicates a smooth transition between the low and high $Q^{2}$ regimes. The results presented here establish a pathway for improved lattice calculations of the subtraction function to deliver improved estimates of the proton-neutron mass difference and the proton charge radius.
The work presented here introduces a number of avenues of future research, particularly by incorporating a consideration of finite volume effects, and by repeating this work for domain wall fermions (see Ref. Fu et al. (2025)). Given the lack of a Wilson term, one would anticipate $\mathcal{O}(a^{2})$ discretisation artefacts for domain wall fermions, as opposed to the $\mathcal{O}(a)$ artefacts present here.
The precision and applicability of our calculation could be further improved by repeating these calculations for pion masses closer to the physical value, and with additional configurations of different lattice spacings. Furthermore, the precision of the approach presented here could be evaluated by extracting $\langle x\rangle$ and $g_{S}$ from fits to the subtraction function, and comparing to the values in Table 1.
@@ Appendix A Details of Lattice Operator Product Expansion
Appendix B Tabulated Subtraction Function Results
Table 3: $\mathcal{O}(a)$-improved subtraction function results for the different configurations, as shown in Fig. 8.
$\kappa=0.120900$
$Q^{2}\,[\text{GeV}^{2}]$ $S_{1}(Q^{2})$
1.10 2.00(32)
1.37 0.984(126)
2.74 0.259(48)
3.56 0.108(79)
4.39 0.074(131)
4.66 0.018(53)
5.48 0.094(63)
6.85 0.047(59)
7.13 -0.006(56)
12.3 0.002(70)
15.9 -0.085(126)
$\kappa=0.122005$
$Q^{2}\,[\text{GeV}^{2}]$ $S_{1}(Q^{2})$
1.44 0.688(71)
1.88 0.596(87)
2.89 0.208(73)
4.91 0.080(69)
7.21 0.136(75)
9.38 0.135(65)
$\kappa=0.122810$
$Q^{2}\,[\text{GeV}^{2}]$ $S_{1}(Q^{2})$
4.96 0.056(98)
6.74 -0.001(75)
9.92 -0.077(103)